solve by the substitution method
8m + n = 11
m - 6n = 32

Answer by Alan3354(69443)

8m + n = 11
m - 6n = 32 --> m = 6n+32
Sub for m into eqn 1
8m + n = 11
8(6n+32) + n = 11
49n + 256 = 11
49n = -245
n = -5
-----------
m = 6n+32 = 6*(-5) + 32
m = 2